OpenStudy (anonymous):

Help, will medal. Find the equation of a line in slope-intercept form that passes through the points (-2, 20) and (3, -10).

OpenStudy (anonymous):

y-20 = -5(x+2) you find slope using m= (y1-y2)/(x1-x2) then just plug in one of the points into the slope intercept form